OpenStudy (anonymous):

let v=(t,-3t) for t in R and let addition and scalar multiplication be the standard operations on vectors. determine whether v is a vector space.

OpenStudy (anonymous):

I think you just have to check if: \[ \mathbf v(x)+\mathbf v(y)=\mathbf v(x+y) \]And if \[ c\mathbf v(x)=\mathbf v(cx) \]

OpenStudy (anonymous):

so one of those will not check out and equal each other? thus it is not a vector space?

OpenStudy (anonymous):

v=(t.-3t) is column vector i just didnt know how to type it as a column vector \[\left(\begin{matrix}t \\ -3t\end{matrix}\right)\]

OpenStudy (anonymous):

It doesn't look like it should be a vector space. v is R (one dimensional) and it has components t and -3t which are dependent on one another. I do not think scaling and adding will be supported

OpenStudy (anonymous):

I'm taking lin aglebra now but have not done proofs, so you should get a second opinion

OpenStudy (anonymous):

Just trying drawing the vector! Its impossible without two lines

OpenStudy (anonymous):

thats what i thought at first that since t is in R and not R^2 that it would not be a vector space since V is t and -3t which that should be in R^2

OpenStudy (anonymous):

I hope I understood the question correctly :-/

OpenStudy (anonymous):

yup thats my logic too!

OpenStudy (anonymous):

what do you mean by drawing the vector and it being impossible without two lines?